OpenStudy (anonymous):

In triangle ABC, what is the measure of A if a = 15, b = 10, and c = 7? 57.12° 72.15° 103.21° 122.88°

OpenStudy (anonymous):

Can you formulate the right law of cosine?

OpenStudy (anonymous):

-2abcos?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

no its a^2=b^2+c^2-2bc(cosa)

OpenStudy (anonymous):

2(10)(7)(cos 15)?

OpenStudy (anonymous):

149

OpenStudy (anonymous):

I just know I did it wrong

OpenStudy (anonymous):

too fast. You need to separate the cos(alpha) part.

OpenStudy (anonymous):

I got 149 for b^2 and c^2

OpenStudy (anonymous):

\[a^2 = b^2 + c^2 + b c \cos(\alpha)\] separate the cos(alpha)

OpenStudy (anonymous):

b^2+c^2 = 149 is true, but does not help you in any way.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

You need to separate the alpha:\[\cos(\alpha) = \frac{a^2-b^2-c^2}{bc}\]\[\alpha = \cos^{-1} \left( \frac{a^2-b^2-c^2}{bc}\right)\] Try it on your own

OpenStudy (anonymous):

\[\cos^{-1} (\frac{ 15^2-10^2-7^2 }{ 10(7)})\]

OpenStudy (anonymous):

i GET 23.56

OpenStudy (anonymous):

@Frouzen

OpenStudy (anonymous):

@Nnesha